Prelab: Chemical Reactions

  1. Name three types of double displacement reactions. How are they defined?

  2. What pushes the double displacement reactions toward completion?

  3. Consider the reaction of aqueous hydrochloric acid and aqueous barium hydroxide. Write balanced molecular, complete ionic, and net ionic equations for this chemical reaction.

  4. What are spectator ions? Identify any spectator ions from the reaction in the question above.

  5. How can one predict whether two aqueous solutions will/will not undergo a double displacement reaction?

  6. A single displacement reaction is a type of redox reaction that follows the general patternWhere the metal, M(s), displaces L+(aq) from the LX(aq) to form the metal, L(s). What is causing this reaction to occur? Explain. M(s) + LX(aq) → MX(aq) + L(s)

  7. Describe the above single displacement reaction in redox terms. What is being oxidized/reduced? What is the reducing agent/oxidizing agent?
Answer #1

M = M+ e- (oxidation)

X + e- = X- (reduction)